# Feature Fusion from Head to Tail: an Extreme Augmenting Strategy for Long-Tailed Visual Recognition

### v1

@@ 4.4 Analysis
  • Note: ${}^{\dagger}$, baselines.
This section validates the efficacy of the fusion strategy proposed in Section 3.3. We compare different sampling strategies including class-balanced sampling (CS), reverse sampling (RS) that is tail-biased, and instance-wise sampling (IS) that is head-biased for the fusing branch. DR uses cRT to fine-tune the classifier with the class-balanced sampling data, which enlarges medium and tail class expands by pushing the decision boundary in the direction of compressing head classes. Additionally, DR uses CE loss that does not consider class margin. These lead to a significant boost in medium and tail classes while reducing the classification accuracy of head classes. RS augments classes of all scales using head class samples with the lowest probability. CS makes all classes augmented with other classes with equal probability. IS allows more features from the medium and tail class samples to be fused with the head features. All sampling strategies adjust the class boundaries towards narrowing the head class distribution. RS has less damage to the performance of head classes while the improvement brought about by IS to the tail and medium classes is more prominent. As the classifier tuning stage cannot change the embedding distribution, the performance on medium and tail classes is further improved but with performance degradation of head classes in the frozen embedding space.
The Impact of Fusion Ratio:
Figure 5: Change of accuracy on different splits w.r.t. $p$.
Figure 5 shows the change of accuracy on different splits with respect to $p$. As the fusion ratio increases, the performance of the medium and tail classes improves while that of the head class decreases, which is consistent with Section 3.4. The class extension of the medium class and the tail class is expanded by eroding that of the head classes through the feature fusion operation. Since $\ddot{w}_{t}^{T}\ddot{f}_{h}>\ddot{w}_{h}^{T}\ddot{f}_{h}$ plays a disproportionate role, resulting in excessive head class samples being incorrectly labeled as medium and tail classes When $p$ increases. As depicted in Figure 5, when $p>0.3$, the accuracy on head classes drops dramatically. Notably, even replacing all feature maps with another class still yields a relatively satisfactory accuracy (43.57%), which surpasses CE loss by 4.02%. $p=0.3$ obtains the best overall performance (47.73%) on CIFAR100-LT. By adjusting $p$ based on specific scenarios, one can control the accuracy of classes with different scales, which is particularly useful when additional accuracy requirements exist for different scale classes. More visualization experiment results can be found in Supplementary A.
@@ 5 Concluding Remarks
Despite its effectiveness, H2T has an underlying assumption that the representation learning has reached its optimum on the available data. It does not alter the feature distribution and only adjusts the decision boundary. Ultimately, improvements in the tail classes are always accompanied by sacrificing the head class performance. Our future work will focus on obtaining a more reasonable embedding space distribution to overcome this limitation.
Supplementary A Visualization Results w.r.t. Different $p$
Figure S1: Visualization of feature distribution w.r.t. $p$.
@@
This supplementary material includes visualization results for fusion ratio $p$ mentioned in Section 3.3 and Section 4.4. We use CIFAR10-LT for easy visualization. Figure S1 shows that as the value of $p$ increases, the interior of each category becomes saturated first, and after $p$ surpasses 0.4, the samples within each category gradually become more dispersed. These findings suggest that the selection of an appropriate fusion ratio is crucial in balancing the preservation of the interior samples while ensuring that the model is able to capture the tail samples effectively. By gaining a deeper understanding of the relationship between the fusion ratio and the distribution of sample points, we can optimize the performance of the model and ensure its suitability for a range of applications.

### v2

@@ 4.4 Further Analysis
The Impact of Fusion Ratio: Figure 6 shows the change of accuracy on different splits with respect to $p$. As $p$ increases, the performance of medium and tail classes improves while that of head classes decreases, which is consistent with Section 3.4.The class extension of the medium class and the tail class is expanded by eroding that of head classes through H2T fusion. Since $\ddot{w}_{t}^{T}\ddot{f}_{h}>\ddot{w}_{h}^{T}\ddot{f}_{h}$ plays a disproportionate role, resulting in excessive head class samples being incorrectly labeled as medium and tail classes as $p$ increases. If $p>0.3$, the accuracy on head classes drops dramatically. Even replacing all feature maps with another class still yields a relatively satisfactory accuracy (43.57%), surpassing CE loss by 4.02%. $p=0.3$ obtains the best overall performance (47.73%) on CIFAR100-LT. By adjusting $p$ based on specific scenarios, one can control the accuracy of classes with different scales, which is particularly useful when additional accuracy requirements exist for different scale classes. More visualization results for the fused embedding space and the selection strategy of fused features can be found in Section E in the Appendix.
@@ 5 Concluding Remarks
Despite its effectiveness, H2T has an underlying assumption that representation learning has reached its optimum on the available data. It only adjusts the decision boundary without altering the feature distribution. Ultimately, improvements in the tail classes are always accompanied by sacrificing the head class performance. Our future work will focus on obtaining a more reasonable embedding space distribution to overcome this limitation.
Acknowledgments
This work was supported in parts by NSFC (62306181, U21B2023, 62376233), DEGP Innovation Team (2022KCX TD025), NSFC-RGC (N_HKBU214/21), RGC (12201321, 12202622), and Shenzhen Science and Technology Program (KQTD20210811090044003, RCJC20200714114435012).
Appendix: Feature Fusion from Head to Tail for Long-Tailed Visual Recognition
Appendix A Key Execution Code
H2T is easy to implement and can be seamlessly integrated with various other methods. Below, we outline the key codes of the algorithm:
Listing 1 PyTorch implementation of H2T
1def H2T(x1, x2, rho = 0.3):
2    if type(x1).__name__ == ’Tensor’:
3        fea_num = x1.shape[1]
4        index = torch.randperm(fea_num).cuda()
5        slt_num = int(rho*fea_num)
6        index = index[:slt_num]
7        x1[:,index,:,:] = x2[:,index,:,:]
8    else:
9        for i in range(len(x1)):
10            fea_num = x1[i].shape[1]
11            index = torch.randperm(fea_num).cuda()
12            slt_num = int(rho*fea_num)
13            index = index[:slt_num]
14            x1[i][:,index,:,:] = x2[i][:,index,:,:]
15    return x1
Appendix B Comparison Results on CIFAR-10-LT
In Section 4.3 of the main paper, we show the results on most of the benchmark datasets except CIRAT10-LT. Table 6 presents the supplementary results on CIFAR10-LT w.r.t Top-1 classification accuracy (%). The improvement of applying H2T to the vanilla training of DR with CE loss (DR+H2T) is more clear. The biggest improvement exceeds 5% on CIFAR10-LT with $\rho=200$. The improvement in GCL+H2T is less compared with other methods because GCL uses the distance in cosine form. This result is consistent with the experimental results on Places-LT (refer to Table 4 of the main paper), namely the H2T improves RIDE with cosine classifier less than with linear classifier.
Table 6: Comparison results on CIFAR10-LT.
Imbalance Ratio 200 100 50
Single Model
CE loss 67.24 70.70 74.81
LDAM-DRW (Cao et al. 2019) 73.52 77.03 81.03
DR+MU (Kang et al. 2020) 73.06 79.15 84.21
MisLAS (Zhong et al. 2021) 77.17 82.12 85.55
${}^{\star}$ (Liu, Li, and Sun 2022) - 81.00 86.60
GCL (Li, Cheung, and Lu 2022) 78.56 82.06 85.36
[…]
MisLAS+H2T 78.05 ($\uparrow$ 0.88) 83.23 ($\uparrow$ 1.11) 86.72 ($\uparrow$ 1.17)
GCL+H2T 79.42 ($\uparrow$ 0.86) 82.40 ($\uparrow$ 0.34) 85.41 ($\uparrow$ 0.05)
Multi-Expert Model
BBN (Zhou et al. 2020) 73.47 79.82 81.18
${}^{\star}$ (Cai, Wang, and Hwang 2021) - 81.20 84.30
ResLT (Cui et al. 2023) 75.29 80.58 84.07
[…]
  • Note: The backbone is ResNet-32. $\star$ represents the results quoted from the corresponding papers. The best and the second-best results are shown in underline bold and bold, respectively. The text in blue indicates the performance improvement compared to the original method w.o. H2T.
Appendix C Selection of fused features
Table 7: Selection of fused features
[…]
47.22 47.23 47.41 47.73
Random 30%_2 Random %30_3 Random 30%_4 Random 30%_5
47.75 47.80 47.71 47.66
The selected fused features are essential. We conduct an experiment to show the effect of the selection strategy of fused features. The dataset is CIRAR100-LT and the imbalance ratio is 100. Table 7 shows the Top-1 classification accuracy (%) of different selections. Two approaches are validated: sequential selection and random selection. Sequential selection refers to replacing a certain proportion of feature maps directly in accordance with the original order of feature maps. We select the first 30%, meddle 30%, and the last 30% of the feature maps to be replaced. Random selection refers to randomly selecting a certain proportion of feature maps to replace. In the experiment, the fused ratio is 30% and we verified 5 random seeds (Random 30%_1 to Random 30%_5). We can observe that random selection generally outperforms sequential selection. Different fused features may introduce variations, underscoring the need for a thoughtful selection strategy. This aspect remains a topic for our future work.
Figure 7: Decision boundary Comparison w.o. and w. H2T (all classes).
Figure 8: Decision boundary Comparison w.o. and w. H2T (Class 0 and 6).
Figure 9: Decision boundary Comparison w.o. and w. H2T (Class 0 and 7).
Refer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to caption
Figure 7: Decision boundary Comparison w.o. and w. H2T (all classes).
Figure 8: Decision boundary Comparison w.o. and w. H2T (Class 0 and 6).
Figure 9: Decision boundary Comparison w.o. and w. H2T (Class 0 and 7).
Figure 10: Decision boundary Comparison w.o. and w. H2T (Class 0 and 9).
Appendix D Visualization Results of decision Boundaries
This section offers additional visualization comparisons pertaining to the decision boundary, serving as a supplementary analysis to Figure 5 presented in Section 4.4 of the main paper. CIFAR10-LT, with an imbalance ratio of 100, is employed for the purpose of facilitating easy visualization. Classes 0-3 are head classes and Classes 6-9 are tail classes. Figure 10 shows the tSNE of embedding space and the decision boundaries. We can see that H2T changes the embedding space of head and tail classes more. Figures 10-10 further visualize the decision boundary between head and each tail class. Compared with DR, H2T fills the margin between classes and adjusts the decision boundary to a more reasonable position in the obtained embedding space. For example, as shown in 10, H2T not only expands the tail class (Class 7) space but also rectifies the head class space (Class 0). While this decision boundary enhances the classification effect, it remains suboptimal. This presents an intriguing problem worth exploring in future research.
Appendix E Visualization Results w.r.t. $p$
@@ Appendix D Visualization Results of decision Boundaries
This section includes visualization results for fusion ratio $p$ mentioned in Section 3.3 of the main paper. We use CIFAR10-LT for easy visualization. Figure 11 shows that as the value of $p$ increases, the interior of each category becomes saturated first, and after $p$ surpasses 0.4, the samples within each category gradually become more dispersed. These findings suggest that the selection of an appropriate fusion ratio is crucial in balancing the preservation of the interior samples while ensuring that the model is able to capture the tail samples effectively. By gaining a deeper understanding of the relationship between the fusion ratio and the distribution of sample points, we can optimize the performance of the model and ensure its suitability for a range of applications.
(Figure 11 is located on the next page.)
Figure 11: Visualization of feature distribution w.r.t. $p$.
Appendix F Comparison with Existing Methods
$\bullet$ Clarify H2T. H2T constructs virtual tail features to simulate potential sample variations by grafting part of head features into the tail, enabling tail samples to populate the class boundary, as depicted in Fig. 1. It enhances generalization by explicitly introducing perturbations.
$\bullet$ Comparison with influence-balanced loss (IB) (Park et al. 2021). Reducing decision boundary bias is a common goal in long-tail learning. Different from H2T, IB implicitly addresses this issue from an optimization perspective. The gradients in stage II of IB is $\lambda_{k}\cdot\left({\nabla\mathcal{L}_{k}}/{\|\nabla\mathcal{L}_{k}\|}\right)$. Essentially, IB directly specifies the gradient descent step size to promote a more substantial gradient descent for tail classes to finetune the classifier.
$\bullet$ Comparison with CMO (Park et al. 2022). CMO allows the transformation of discrete sample values into continuous ones by mixup inputs and labels using CutMix. No new samples are created essentially. H2T constructs new virtual tail features where labels are not involved in fusion, which can simulate samples that are absent from the training set.
$\bullet$ Comparison with head-to-tail network (HTTN) (Xiao et al. 2021). HTTN indirectly utilizes head samples to transfer the prototypes. It computes the attention ($\alpha_{zj}$) between tail and head prototypes, then uses $\alpha_{zj}$ and the relationship ($W_{transfer}$) between head prototypes and head classifier to identify tail classifier. In H2T, tail classes directly employ a portion of head features to simulate unseen tail samples and thus populate the class boundary.
Appendix G Additional Rational Analysis
The addition of the inequality obtained from correctly classified tail samples to Eq. 4 does not eliminate the common terms on both sides of the inequality, and consequently, it does not lead to a meaningful result. We consider the inequality relationship obtained by correctly classified tail samples and fused head samples:
[…]
Then, we can derive Eq. 9, which encourages $w_{t}$ to be in closer proximity to tail samples, as depicted in Figure 4.